OpenStudy (anonymous):

How would the graph of y = x^2 differ from the graph of y = –3(x – 2)^2 + 1?

OpenStudy (anonymous):

y = -3(x-2)^2 +1 is a transformed graph of y = x^2 with the following transformations: 1. Vertical stretched by a factor of 3 2. Reflection on the x axis 3. Horizontal translation 2 units right 4. Vertical translation 1 unit up